# Agent: Generating and Evaluating Presentations Beyond Text-to-Slides

### v1

@@ 4.3 Evaluation Metrics
Success Rate (SR) measures the robustness of the generation task by determining the percentage of presentations where all slides are successfully generated.
@@
Perplexity (PPL) measures the likelihood of the language model generating the given sequence. Following Bandyopadhyay et al. (2024), we calculate the average perplexity of slides within a presentation using GPT-2.. A lower perplexity score indicates that the textual content is more fluent.
@@
FID (Heusel et al., 2017) measures the similarity between the generated presentation and the exemplar presentation in the feature space. Due to the limited sample size, we calculate the FID using a 64-dimensional output vector.
PPTEval measures the comprehensive quality of presentations across three dimensions: coherence, content, and design. We employ GPT-4o as the judge model.
4.4 Result & Analysis
@@
Table 2 presents the performance comparison between Agent and baseline methods, revealing that:
PPTAgent Enhances LLMs’ Presentation Generation Capabilities
As demonstrated in Table 2, our approach empowers LLMs to produce well-rounded presentations with a remarkable success rate, achieving success rate for both Qwen2.5+Qwen2-VL and GPT-4o+GPT-4o. This is a significant improvement compared to the highest accuracy of 10% for session-based template editing tasks as reported in Guo et al. (2023). This improvement can be attributed to three main factors: 1) Agent concentrates on content modifying, thereby avoiding intricate stying operations. 2) Our streamlined API design allows LLMs to execute tasks with ease. 3) The code interaction module enhances LLMs’ comprehension of slides and offers opportunities for self-correction, enabling them to generate accurate actions robustly. Moreover, detailed performance of Qwen2.5+Qwen2-VL across various domains, as illustrated in Table 3, underscores the robustness of our approach.
PPTAgent Significantly Improves Overall Presentation Quality
By adopting an Edit-based paradigm, Agent allows elements within the presentation to inherit well-designed styling attributes from existing presentations. When using GPT-4o, experimental results demonstrate comprehensive improvements over the baseline. We significantly surpass the baseline method in the design dimension under Eval (3.24 vs 2.33), as the presentations generated by the baseline method lack basic design efforts. Furthermore, we achieved substantial enhancements in coherence (4.39 vs 3.28) and content (3.25 vs 2.98) dimensions, as the semantic information extracted during the Presentation Analysis stage effectively guided the LLMs.
Open-Source LLMs Rival GPT-4o in Performance
GPT-4o consistently demonstrates outstanding performance across various evaluation metrics, highlighting its advanced capabilities. While Qwen2-VL exhibits limitations in linguistic proficiency due to the trade-offs from multimodal post-training, GPT-4o maintains a clear advantage in handling language tasks. However, the introduction of Qwen2.5 successfully mitigates these linguistic deficiencies, bringing its performance on par with GPT-4o, and achieving the best performance. This underscores the significant potential of open-source LLMs as competitive and highly capable presentation agents.
@@ 4.5 Ablation Study
To better understand the impact of each component in our proposed method, we performed ablation studies using four different configurations. Specifically, we evaluated the method by: (1) randomly selecting a slide as the edit target (w/o Outline), (2) omitting structural information during outline generation (w/o Structure), (3) replacing the slide representation with the method described in Guo et al. (2023) (w/o CodeRender), and (4) removing guidance from the slide schema (w/o Schema). These configurations were tested using the Qwen2.5+Qwen2-VL.
Code Representation Enhances LLMs’ Comprehension
As shown in Table 2, the removal of the Code Render component leads to a significant drop in the model’s success rate (SR) from 95.0 to 74.6. This underscores the critical role of code representation in leveraging LLMs’ coding capabilities to improve their overall comprehension.
Presentation Analysis is Essential for Generating Targeted Presentations
The removal of the outline and structural information significantly degrades coherence (from 4.48 to 3.36/3.45), underscoring their crucial role in maintaining logical flow. Furthermore, the absence of slide schema hinders LLMs from generating targeted content effectively, resulting in a drop in success rate from 95.0 to 78.8.
4.6 Error Analysis
@@
Figure 4 illustrates the number of iterations required to generate a slide using different models. Although GPT-4o exhibits superior self-correction capabilities compared to Qwen2.5, Qwen2.5 encounters fewer errors in the first iteration (Iter-0). Additionally, we observed that Qwen2-VL experiences errors more frequently and has poorer self-correction capabilities, likely due to its multimodal post-training (Wang et al., 2024a). Ultimately, all three models successfully corrected more than half of the errors, demonstrating that our iterative self-correction mechanism effectively ensures the success of the generation process.
| Corelation | Content | Design | Coherence | Avg. |
@@
4.7 Effectiveness of PPTEval
Human Agreement Evaluation
Despite Chen et al. (2024a) have highlighted the impressive human-like discernment of LLMs in various generation tasks. However, it remains crucial to assess the correlation between LLM evaluations and human evaluations in the context of presentations. This necessity arises from findings by Laskar et al. (2024), which indicate that LLMs may not be adequate evaluators for complex tasks. Table 4 shows the correlation of ratings between humans and LLMs. The average Pearson correlation of 0.71 exceeds the scores of other evaluation methods (Kwan et al., 2024), indicating that Eval aligns well with human preferences.
Moreover, the heatmap in Figure 5 reveals the limitations of existing metrics when compared with the Content and Design dimensions of Eval. In our experiments, we observed that PPL predominantly captures text fluency and is susceptible to the fragmented nature of slide text, leading to ineffective measurements with frequent outliers. Similarly, FID merely quantifies stylistic similarity to reference presentations rather than design quality, as conformity to reference styles does not necessarily indicate superior design. These findings underscore the necessity of Eval for comprehensive and effective presentation evaluation.
5 Related Works
Automated Presentation Generation
Recent proposed methods for slide generation can be categorized into rule-based and template-based based on how they handle element placement. Rule- based methods, such as those proposed by Mondal et al. (2024) and Li et al. (2021), often focus on enhancing textual content but neglect the visual-centric nature of presentations, leading to outputs that lack engagement. Template-based methods, including Cachola et al. (2024) and industrial solutions like Tongyi, rely on pre-designed templates to create visually appealing presentations. However, their dependence on extensive manual effort for template annotation significantly limits scalability and flexibility.
LLM Agent
Numerous studies (Li et al., 2024; Deng et al., 2024; Wang et al., 2024c) have explored the potential of LLMs to act as agents assisting humans in a wide array of tasks. For example, Zheng et al. (2024); Wang et al. (2024b) demonstrate the capability of LLMs to accomplish tasks by generating executable actions and correcting errors based on feedback. Furthermore, Guo et al. (2023) introduces an evaluation system that assesses the ability of LLMs to perform multi-turn, multimodal slide editing tasks using APIs, which inspired the use of LLMs for complex tasks as proposed in this study.
LLM as a Judge
LLMs have demonstrated strong capabilities in instruction following and context perception, leading to their widespread use as judges (Liu et al., 2023; Zheng et al., 2023). Further research by Zhuge et al. (2024) enhanced LLMs’ abilities through external modules and functions, while Chen et al. (2024a) validated the feasibility of using multimodal large language models(MLLMs) as judges. Additionally, Kwan et al. (2024) introduced a multi-dimensional evaluation framework for multi-turn conversations, which inspired the development of our proposed Eval.
@@ 6 Conclusion
In this paper, we introduced Agent, which conceptualizes presentation generation as a two-stage presentation editing task completed through the abilities of LLMs to understand and generate code. This approach leveraged the textual feature and layout patterns to organize slides into different functional groups. Our experiments across data from multiple domains have demonstrated the superiority of our method. Moreover, our proposed Eval ensured the assessability of presentations. This research provides a new paradigm for generating slides under unsupervised conditions and offers fresh insights for future work in presentation generation.
7 Limitations
While our method demonstrates its capability to produce high-quality presentations, there remain inherent challenges that impact its universal applicability. For instance, achieving a success rate of over 95% on our dataset is impressive, but not absolute, thus might limit its application. Moreover, parsing slides with intricate nested group shapes often proves to be a bottleneck, leading to less consistent results. Additionally, although Agent shows noticeable improvements in layout optimization over prior approaches, it still falls short of exploiting the full potential of visual cues for refining stylistic consistency. This often manifests in design flaws, such as overlapping elements, undermining the visual harmony of the generated slides. Addressing these limitations calls for future enhancements that integrate visual information into the generation process.
8 Ethical Considerations
In the construction of Zenodo10K, we utilized the publicly available API to scrape data while strictly adhering to the licensing terms associated with each artifact. Specifically, artifacts that were not permitted for modification or commercial use under their respective licenses were filtered out to ensure compliance with intellectual property rights. Additionally, all annotation personnel involved in the project were compensated at rates exceeding the minimum wage in their respective cities, reflecting our commitment to fair labor practices and ethical standards throughout the dataset’s development process.
Appendix A Data Sampling
To maintain a reasonable cost, we selected presentations ranging from 12 to 64 pages and documents with text lengths from 2,048 to 20,480 characters.
@@ Appendix B Details of PPTEval
Through a Shanghai-based crowdsourcing platform, we recruited four graduate students to evaluate 50 randomly selected presentations from Zenodo10K, along with 100 presentations generated by the baseline method and our approach, respectively. The evaluations were conducted across three dimensions, as proposed by Eval, based on the same scoring criteria listed in Appendix E along with converted slide images. Moreover, we listed some scoring examples in Figure 6 and detailed performance of Qwen2.5+Qwen2-VL across various domains in Table 3.
Appendix C Layout Analysis
We present our hierarchical clustering algorithm for layout analysis in Algorithm 1, where slides are grouped into clusters using a similarity threshold of 0.65. To minimize clustering interference, we replace the text and images in the slides with placeholders beforehand. Moreover, examples of the extracted slide clusters are provided in Figure 7.
Appendix D Code Interaction
Appendix E Prompts
E.1 Prompts for Presentation Analysis
E.2 Prompts for Presentation Generation
E.3 Prompts for PPTEval
| Function Name | Description |
| del_span | Deletes a specific span. |
| del_image | Deletes an image element. |
| clone_paragraph | Creates a duplicate of an existing paragraph. |
| replace_span | Replaces the content of a specific span. |
| replace_image | Replaces an image with a new image. |

### v3

@@ 4.4 Evaluation Metrics
Success Rate (SR) evaluates the robustness of presentation generation (Wu et al., 2024), calculated as the percentage of successfully completed tasks. For PPTAgent, success requires the generation of all slides without execution errors after self-correction. For KCTV, success is determined by the successful compilation of the generated LaTeX file. DocPres is excluded from this evaluation due to its deterministic rule-based conversion.
@@
Perplexity (PPL) measures the likelihood of the model generating the given sequence. Using Llama-3-8B (Dubey et al., 2024), we calculate the average perplexity across all slides in a presentation. Lower perplexity scores indicate higher textual fluency (Bandyopadhyay et al., 2024).
Rouge-L (Lin, 2004) evaluates textual similarity by measuring the longest common subsequence between generated and reference texts. We report the F1 score to balance precision and recall.
@@
FID (Heusel et al., 2017) measures the similarity between the generated presentation and the reference presentation in the feature space. Due to the limited sample size, we calculate the FID using a 64-dimensional output vector.
PPTEval employs GPT-4o as the judging model to evaluate presentation quality across three dimensions: content, design, and coherence. We compute content and design scores by averaging across slides, while coherence is assessed at the presentation level.
4.5 Overall Result
@@
Table 3 presents the performance comparison between PPTAgent and baselines, revealing that:
Setting | SR(%) | Content | Design | Coherence | Avg. |
PPTAgent | 95.0 | 3.28 | 3.27 | 4.48 | 3.67 |
w/o Outline | 91.0 | 3.24 | 3.30 | 3.36 | 3.30 |
w/o Schema | 78.8 | 3.08 | 3.23 | 4.04 | 3.45 |
w/o Structure | 92.2 | 3.28 | 3.25 | 3.45 | 3.32 |
w/o CodeRender | 74.6 | 3.27 | 3.34 | 4.38 | 3.66 |
PPTAgent Significantly Improves Overall Presentation Quality.
PPTAgent demonstrates statistically significant performance improvements over baseline methods across all three dimensions of PPTEval. Compared to the rule-based baseline (DocPres), PPTAgent exhibits substantial improvements in both the design and content dimensions (3.34 vs. 2.37, +40.9%; 3.34 vs. 2.98, +12.1%), as presentations generated by the DocPres method show minimal design effort. In comparison with the template-based baseline (KCTV), PPTAgent also achieves notable improvements in both design and content (3.34 vs. 2.95, +13.2%; 3.28 vs. 2.55, +28.6%), underscoring the efficacy of the edit-based paradigm. Most notably, PPTAgent shows a significant enhancement in the coherence dimension (4.48 vs. 3.57, +25.5% for DocPres; 4.48 vs. 3.28, +36.6% for KCTV). This improvement can be attributed to PPTAgent ’s comprehensive analysis of the structural role of slides.
PPTAgent Exhibits Robust Generation Performance.
Our approach empowers LLMs to produce well-rounded presentations with remarkable success rate, achieving success rate for both Qwen2.5+Qwen2-VL and GPT-4o+GPT-4o, which is a significant improvement compared to KCTV (97.8% vs. 88.0%). Moreover, detailed performance of Qwen2.5+Qwen2-VL across various domains is illustrated in Table 8, underscoring the versatility and robustness of our approach.
PPTEval Demonstrates Superior Evaluation Capability.
Traditional metrics like PPL and ROUGE-L demonstrate inconsistent evaluation trends compared to PPTEval. For instance, KCTV achieves a high ROUGE-L (16.76) but a low content score (2.55), while our method shows the opposite trend with ROUGE-L (14.25) and content score (3.28). Moreover, we observe that ROUGE score overemphasizes textual alignment with source documents, potentially compromising the expressiveness of presentations. Most importantly, PPTEval advances beyond existing metrics through its dual capability of reference-free design assessment and holistic evaluation of presentation coherence. Further agreement evaluation is shown in Section 5.5.
5 Analysis
@@ 5.1 Ablation Study
We conducted ablation studies across four settings: (1) randomly selecting a slide as the reference (w/o Outline), (2) omitting structural slides during outline generation (w/o Structure), (3) replacing the slide representation with the method proposed by Guo et al. (2023) (w/o CodeRender), and (4) removing guidance from the content schema (w/o Schema). All experiments were conducted using the Qwen2.5+Qwen2-VL configuration.
As demonstrated in Table 4, our experiments reveal two key findings: 1) The HTML-based representation significantly reduces interaction complexity, evidenced by the substantial decrease in success rate from 95.0% to 74.6% when removing the Code Render component. 2) The presentation analysis is crucial for generation quality, as removing the outline and structural slides significantly degrades coherence (from 4.48 to 3.36/3.45) and eliminating the slide schema reduces the success rate from 95.0% to 78.8%.
5.2 Case Study
We present representative examples of presentations generated under different configurations in Figure 5. PPTAgent demonstrates superior presentation quality across multiple dimensions. First, it effectively incorporates visual elements with contextually appropriate image placements, while maintaining concise and well-structured slide content. Second, it exhibits diversity in generating visually engaging slides under diverse references. In contrast, baseline methods (DocPres and KCTV) produce predominantly text-based slides with limited visual variation, constrained by their rule-based or template-based paradigms.
5.3 Score Distribution
We further investigated the score distribution of generated presentations to compare the performance characteristics across methods, as shown in Figure 4. Constrained by their rule-based or template-based paradigms, baseline methods exhibit limited diversity in both content and design dimensions, with scores predominantly concentrated at levels 2 and 3. In contrast, PPTAgent demonstrates a more dispersed score distribution, with the majority of presentations (>80%) achieving scores of 3 or higher in these dimensions. Furthermore, due to PPTAgent’s comprehensive consideration of structural slides, it achieves notably superior coherence scores, with over 80% of the presentations receiving scores above 4.
5.4 Effectiveness of Self-Correction
@@
Figure 6 illustrates the number of iterations required to generate a slide using different language models. Although GPT-4o exhibits superior self-correction capabilities compared to Qwen2.5, Qwen2.5 encounters fewer errors in the first generation. Additionally, we observed that Qwen2-VL experiences errors more frequently and has poorer self-correction capabilities, likely due to its multimodal post-training (Wang et al., 2024a). Ultimately, all three models successfully corrected more than half of the errors, demonstrating that our iterative self-correction mechanism effectively ensures the success of the generation process.
5.5 Agreement Evaluation
PPTEval with Human Preferences
Despite Chen et al. (2024a) have highlighted the impressive human-like discernment of LLMs in various generation tasks. However, it remains crucial to assess the correlation between LLM evaluations and human evaluations in the context of presentations. This necessity arises from findings by Laskar et al. (2024), which indicate that LLMs may not be adequate evaluators for complex tasks. Table 5 shows the correlation of ratings between humans and LLMs. The average Pearson correlation of 0.71 exceeds the scores of other evaluation methods (Kwan et al., 2024), indicating that PPTEval aligns well with human preferences.
PPTEval with Existing Metrics
We analyzed the relationships between PPTEval’s content and design dimensions and existing metrics through Pearson correlation analysis, as shown in Figure 7. The Pearson correlation coefficients reveal that current metrics are ineffective for presentation evaluation. Specifically, PPL primarily measures text fluency but performs poorly on slide content due to its inherent fragmented nature, frequently producing outlier measurements. Similarly, while ROUGE-L and FID quantify similarity to reference text and presentations respectively, these metrics inadequately assess content and design quality, as high conformity to references does not guarantee presentation effectiveness. These weak correlations highlight the necessity of PPTEval for robust and comprehensive presentation evaluation that considers both content quality and design effectiveness.
6 Related Works
Automated Presentation Generation
Recent proposed methods for slide generation can be categorized into rule-based and template-based based on how they handle element placement and styling. Rule-based methods, such as those proposed by Mondal et al. (2024) and Bandyopadhyay et al. (2024), often focus on enhancing textual content but neglect the visual-centric nature of presentations, leading to outputs that lack engagement. Template-based methods, including Cachola et al. (2024) and industrial solutions like Tongyi, rely on predefined templates to create visually appealing presentations. However, their dependence on extensive manual effort for template annotation significantly limits scalability and flexibility.
LLM Agent
Numerous studies (Li et al., 2024; Deng et al., 2024; Tang et al., 2025) have explored the potential of LLMs to act as agents assisting humans in a wide array of tasks. For example, Wang et al. (2024b) demonstrate the capability of LLMs to accomplish tasks by generating executable actions. Furthermore, Guo et al. (2023) demonstrated the potential of LLMs in automating presentation-related tasks through API integration.
LLM as a Judge
LLMs have exhibited strong capabilities in instruction following and context perception, which has led to their widespread adoption as judges (Liu et al., 2023; Zheng et al., 2023). Chen et al. (2024a) demonstrated the feasibility of using MLLMs as judges, while Kwan et al. (2024) proposed a multi-dimensional evaluation framework. Additionally, Ge et al. (2025) investigated the use of LLMs for assessing single-slide quality. However, they did not evaluate presentation quality from a holistic perspective.
Correlation | Content | Design | Coherence | Avg. |
@@ 7 Conclusion
In this paper, we introduce PPTAgent, which conceptualizes presentation generation as a two-stage presentation editing task completed through LLMs’ abilities to understand and generate code. Moreover, we propose PPTEval to provide quantitative metrics for assessing presentation quality. Our experiments across data from multiple domains have demonstrated the superiority of our method. This research provides a new paradigm for generating slides under unsupervised conditions and offers insights for future work in presentation generation.
Limitations
While PPTAgent demonstrates promising capabilities in presentation generation, several limitations remain. First, despite achieving a high success rate (>95%) on our dataset, the model occasionally fails to generate presentations, which could limit its reliability. Second, although we can provide high-quality preprocessed presentations as references, the quality of generated presentations is still influenced by the input reference presentation, which may lead to suboptimal outputs. Third, although PPTAgent shows improvements in layout optimization compared to prior approaches, it does not fully utilize visual information to refine the slide design. This manifests in occasional design flaws, such as overlapping elements, which can compromise the readability of generated slides. Future work should focus on enhancing the robustness, reducing reference dependency, and better incorporating visual information into the generation process.
Ethical Considerations
In the construction of Zenodo10K, we utilized the publicly available API to scrape data while strictly adhering to the licensing terms associated with each artifact. Specifically, artifacts that were not permitted for modification or commercial use under their respective licenses were filtered out to ensure compliance with intellectual property rights. Additionally, all annotation personnel involved in the project were compensated at rates exceeding the minimum wage in their respective cities, reflecting our commitment to fair labor practices and ethical standards.
Appendix A Data Preprocessing
To maintain a reasonable cost, we selected presentations ranging from 12 to 64 pages and documents with text lengths from 2,048 to 20,480 characters. We extracted both textual and visual content from the source documents using VikParuchuri (2023). The extracted text was then organized into sections. For visual content, we generated image captions to assist in relevant image selection through textual descriptions. To minimize redundancy, we identified and removed duplicate images if their image embeddings had a cosine similarity score exceeding 0.85. For slide-level deduplication, we removed individual slides if their text embeddings had a cosine similarity score above 0.8 compared to the preceding slide, as suggested by Fu et al. (2022).
@@ Appendix B Details of PPTEval
We recruited four graduate students through a Shanghai-based crowdsourcing platform to evaluate a total of 250 presentations: 50 randomly selected from Zenodo10K representing real-world presentations, along with two sets of 100 presentations generated by the baseline method and our approach respectively. Following the evaluation framework proposed by PPTEval, assessments were conducted across three dimensions using the scoring criteria detailed in Appendix F. Evaluators were provided with converted slide images, scored them individually, and then discussed the results to reach a consensus on the final scores.
Moreover, We measured inter-rater agreement using Fleiss’ Kappa, with an average score of 0.59 across three dimensions (0.61, 0.61, 0.54 for Content, Design, and Coherence, respectively) indicating satisfactory agreement (Kwan et al., 2024) among evaluators. Representative scoring examples are shown in Figure 8.
We provided detailed illustration as below:
Content:
Design:
Coherence:
Appendix C Detailed Performance of PPTAgent
We present a detailed performance analysis of Qwen2.5LM+Qwen2-VLVM across various domains in Table 8. Additionally, Table 7 and 6 show the success rate-weighted performance, where failed generations receive a PPTEval score of 0, demonstrating that a lower success rate significantly impacts the overall effectiveness of the method.
As demonstrated in Table 6. GPT-4o consistently demonstrates outstanding performance across various evaluation metrics, highlighting its advanced capabilities. While Qwen2-VL exhibits limitations in linguistic proficiency due to the trade-offs from multimodal post-training, GPT-4o maintains a clear advantage in handling language tasks. However, the introduction of Qwen2.5 successfully mitigates these linguistic deficiencies, bringing its performance on par with GPT-4o, and achieving the best performance. This underscores the significant potential of open-source LLMs as competitive and highly capable presentation agents.
Appendix D Slide Clustering
We present our hierarchical clustering algorithm for layout analysis in Algorithm 1, where slides are grouped into clusters using a similarity threshold of 0.65. To focus exclusively on layout patterns and minimize interference from specific content, we preprocess the slides by replacing text content with a placeholder character (“a”) and substituting image elements with solid-color backgrounds. Then, we compute the similarity matrix using cosine similarity based on the ViT embeddings of converted slide images between each slide pair. Figure 9 illustrates representative examples from the resulting slide clusters.
Appendix E Code Interaction
Appendix F Prompts
F.1 Prompts for Presentation Analysis
F.2 Prompts for Presentation Generation
F.3 Prompts for PPTEval
Configuration | Existing Metrics | PPTEval | |||||||
Language Model | Vision Model | SR(%) | PPL | ROUGE-L | FID | Content | Design | Coherence | Avg. |
DocPres (rule-based) | |||||||||
GPT-4o | – | – | 76.42 | 13.28 | – | 2.98 | 2.33 | 3.24 | 2.85 |
Qwen2.5 | – | – | 100.4 | 13.09 | – | 2.96 | 2.37 | 3.28 | 2.87 |
KCTV (template-based) | |||||||||
GPT-4o | – | 80.0 | 68.48 | 10.27 | – | 1.99 | 2.35 | 2.85 | 2.40 |
Qwen2.5 | – | 88.0 | 41.41 | 16.76 | – | 2.24 | 2.59 | 2.95 | 2.59 |
PPTAgent (ours) | |||||||||
GPT-4o | GPT-4o | 97.8 | 721.54 | 10.17 | 7.48 | 3.17 | 3.16 | 4.20 | 3.54 |
Qwen2-VL | Qwen2-VL | 43.0 | 265.08 | 13.03 | 7.32 | 1.34 | 1.43 | 1.75 | 1.50 |
Qwen2.5 | Qwen2-VL | 95.0 | 496.62 | 14.25 | 6.20 | 3.11 | 3.10 | 4.25 | 3.48 |
Setting | SR(%) | Content | Design | Coherence | Avg. |
PPTAgent | 95.0 | 3.11 | 3.10 | 4.25 | 3.48 |
w/o Outline | 91.0 | 2.94 | 3.00 | 3.05 | 3.00 |
w/o Schema | 78.8 | 2.42 | 2.54 | 3.18 | 2.71 |
w/o Structure | 92.2 | 3.02 | 2.99 | 3.18 | 3.06 |
w/o CodeRender | 74.6 | 2.43 | 2.49 | 3.26 | 2.73 |